Find all possible Taylor's series and Laurent series expansions of f(z)= (2z - 3)/((z - 2)(z - 1)) about z=0?


1
Expert's answer
2022-01-19T16:23:59-0500

Resolving into partial fraction.

"f(z)= \frac{1}{z-2}+\frac{1}{z-1}\\
\text{for }\frac{1}{z-1},\text{in not negative power of z}\\
\frac{1}{z-1} = -(\frac{1}{1-z}) = -\sum_{n=0}^\infty z^n~~~valid ~if~|z|<1\\
\text{for }\frac{1}{z-1},\text{in negative power of z}\\
\frac{1}{z-1} =\frac{1}{z(1-z^{-1})} = \sum_{n=0}^\infty \frac{1}{z^{n+1}}~~valid ~if~|z|>1\\

\text{for }\frac{1}{z-2},\text{in not negative power of z}\\

\frac{1}{z-2} =\frac{1}{z(1-\frac{1}{2}z)} = \sum_{n=0}^\infty \frac{2^n}{z^{n+1}}~~valid ~if~|z|>2\\

\frac{1}{z-2} =-[\frac{1}{2(1-\frac{1}{2}z)}] = -\sum_{n=0}^\infty \frac{z^n}{2^{n+1}}~~valid ~if~|z|<2\\
\\

f(z)=-\sum_{n=0}^\infty (1+\frac{1}{2^{n+1}}z^n)~~valid ~if~|z|<1\\

f(z)=-\sum_{n=0}^\infty ({2^{n+1}}) \frac{1}{z^n}~~valid ~if~|z|<1\\"